22. Given the cash flows below, answer the following question (provide your answer and show your work in your uploaded document) (maximum points: 3.33)

                                    Date                Cash Flow       Balance

                                    7/1                      Begin            2,150

                                    7/31                    ------              2,280

                                    8/1                      $180             2,460

                                    8/31                    ------              2,561

                                    9/1                      $100             2,661

                                    9/30                 End               2,500

Calculate the average monthly rate of return using the money (dollar)-weighted method over the three months provided here (round your answer to 3 decimals in the form 0.000) Show your work in the uploaded document:
Answer #1
Date balances money weighted return
7/1 2150
7/31 2280 =2150/2280=0.943
8/1 2460
8/31 2561 =2460/2561=0.960
9/1 2661
9/30 2500 =2661/2500=1.064

Average monthly retrun for 3 months is = ( 0.943+0.960+1.064 )/3

= 0.989